Find initial solution for the following transportation problem using Least Cost method. D1 D2 D3 D4 Availability S1 19 30 50 10 7 S2 70 30 40 60 9 S3 40 8 70 20 18 Requirements 5 8 7 14 


1
Expert's answer
2022-02-09T11:55:24-0500

Solution:

TOTAL number of supply constraints : 3

TOTAL number of demand constraints : 4

Problem Table is



The minimum total transportation cost =10×7+70×2+40×7+40×3+8×8+20×7=814


Here, the number of allocated cells = 6 is equal to m + n - 1 = 3 + 4 - 1 = 6

∴ This solution is non-degenerate